# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v1

@@ Acknowledgements.
The authors wish to thank R. Lecamwasam, D. Porter, A. M. Nicholls and B. Shajilal for helpful discussions. The theoretical work was supported by Grant-in-Aid for Scientific Research (Grant No. 22K03507) from the Japan Society for the Promotion of Science and the Iketani Science and Technology Foundation. We also acknowledge support from the Singapore National Science Scholarship, Agency for Science Technology and Research and the European Research Council (HERO, Grant No. 810451). The RIXS experiment at the rhenium $L_{3}$ edge was supported by “Advanced Research Infrastructure for Materials and Nanotechnology in Japan (ARIM)” of the Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan (Proposal No. JPMXP1222QS0107) and performed at BL11XU beamline of SPring-8 with the approval of the Japan Synchrotron Radiation Research Institute (JASRI) (Proposal No. 2022B3596). The O $K$-edge RIXS experiments (Proposal No. 20222125) were performed at the ADRESS beamline of the Swiss Light Source at the Paul Scherrer Institut (PSI). The experimental work at PSI is supported by the Swiss National Science Foundation through project nos. 178867 and 207904. T.Y. is funded by a CROSS project of PSI. Y.W. and G.C.W. acknowledge funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement No. 884104 (PSIFELLOW-II-3i program).

### v2

@@ Acknowledgements.
The authors wish to thank R. Lecamwasam, D. Porter, A. M. Nicholls, and B. Shajilal for their helpful discussions. The theoretical work was supported by a Grant-in-Aid for Scientific Research (Grant No. 22K03507) from the Japan Society for the Promotion of Science, the Iketani Science and Technology Foundation, and the Chiba University Open Recruitment for International Exchange Program. We also acknowledge support from the Singapore National Science Scholarship, Agency for Science Technology and Research and the European Research Council (HERO, Grant No. 810451). The RIXS experiment at the rhenium $L_{3}$ edge was supported by “Advanced Research Infrastructure for Materials and Nanotechnology in Japan (ARIM)” of the Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan (Proposal No. JPMXP1222QS0107) and performed at BL11XU beamline of SPring-8 with the approval of the Japan Synchrotron Radiation Research Institute (JASRI) (Proposal No. 2022B3596). The O $K$-edge RIXS experiments (Proposal No. 20222125) were performed at the ADRESS beamline of the Swiss Light Source at the Paul Scherrer Institut (PSI). The experimental work at PSI is supported by the Swiss National Science Foundation through project nos. 178867 and 207904. T.Y. is funded by a CROSS project of PSI. Y.W. and G.C.W. acknowledge funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement No. 884104 (PSIFELLOW-II-3i program).
Appendix A Computational details
We obtained the vibronic states of an isolated $5d^{1}$ octahedron by numerically diagonalizing the model Hamiltonian consisting of Eqs. (1), (2), and (3) [See Refs. [15, 13]]. To this end, we constructed the Hamiltonian matrix with the basis set of $|\Gamma\gamma\rangle\otimes|n_{z^{2}},n_{x^{2}-y^{2}}\rangle$ ($n_{z^{2}},n_{x^{2}-y^{2}}=0,1,2,...$, and $n_{z^{2}}+n_{x^{2}-y^{2}}\leq 15$). Here, $|n_{z^{2}},n_{x^{2}-y^{2}}\rangle$ stands for the eigenstates of the two-dimensional harmonic oscillator in Eq. (2).
To include the mean-field effect into the vibronic states, we numerically diagonalized the above Hamiltonian with Eq. (6). We determined $F_{z^{2}}=3.3$ meV so that the expectation value of $\hat{q}_{z^{2}}$ corresponds to the experimental structure at 17 K [21].
We consider the RIXS processes involving the transitions between the Re $2p$/O $1s$ to the $5d$ $t_{2g}$ orbitals. For the calculations of the RIXS intensities, we employed the Kramers-Heisenberg formula [27, 28].
$I(\omega_{k^{\prime}})$ $\propto\sum_{\nu}\rho_{\nu}\frac{\omega_{k^{\prime}}}{\omega_{k}}\left|\sum_{\mu}\frac{\langle\nu^{\prime}|\hat{P}_{\bm{k}^{\prime}\alpha^{\prime}}|\mu\rangle\langle\mu|\hat{P}_{\bm{k}\alpha}|\nu\rangle}{E_{\nu}+\hslash\omega_{k}-E_{\mu}+i\Gamma_{\mu}}\right|^{2}$
×$\times\delta\left(E_{\nu^{\prime}}+\hslash\omega_{k^{\prime}}-E_{\nu}-\hslash\omega_{k}\right),$ (7)
where $\hat{P}_{\bm{k}\alpha}$ is the dipole moments along the polarization vector of the light, and $\rho_{\nu}$ is the canonical distribution for single $5d^{1}$ site. The initial and final states denoted by $\nu,\nu^{\prime}$ in Eq. (7) are the vibronic types, and the intermediate states $\mu$ are $5d^{2}$ multiple type. The broadening of the intermediate states $\Gamma_{\mu}$ is 150 meV [18]. In the Re $L_{3}$-edge RIXS calculations, we ignored the multiplet structure of the intermediate states as in Ref. [13]. In the O $K$-edge RIXS, the transition occurs from the O $1s$ orbital to the Re 5d-like $t_{2g}$ orbitals hybridized with the O $2p$ orbitals. To reproduce the RIXS spectra, we convoluted the intensities (7) with the Lorentzian function: The linewidths are $\Gamma=115$ meV for Re $L_{3}$- and $\Gamma=35$ meV for O $K$-edge RIXS spectra, respectively. We calculated only the inelastic contributions for the O $K$-edge RIXS spectra.
We determined $\lambda$ and $g$ so that the simulated Re $L_{3}$-edge RIXS spectra agree with the experimental spectra [Fig. 4(a)].To find the best parameters, we varied the spin-orbit gap by 5 meV and $g$ by 0.05. The former linearly shifts the position of the $j_{\text{eff}}=1/2$ peak, and $g$ varies the shape as shown in Fig. 2.